A balloon has a volume of 13.7 L at 34.0 ºC. What would be the volume of the balloon at –12.0 ºC?


1
Expert's answer
2021-04-28T06:17:50-0400

From Charles law


"\frac{V_1}{T_1}=\frac{V_2}{T_2}"


"\frac{13.7}{307}= \frac{V_2}{261}"


"V_1=\frac{13.7×261}{307}= 11.65 L"